OpenStudy (anonymous):

sqrt(-64) /(7-6i)-(2-2i)

OpenStudy (anonymous):

anyone know this?

OpenStudy (cwtan):

yes here i come sqrt-64=8i The denominator is (7-6i)-(2-2i)?

OpenStudy (anonymous):

how do i distribute the denominator?

OpenStudy (cwtan):

multiply to it's conjugate

OpenStudy (anonymous):

so the conjugate of both ?

OpenStudy (cwtan):

multiply to (conjugate of denominator)/(conjugate of denominator) Is this clear?

OpenStudy (anonymous):

would it be 14+12i ?

OpenStudy (cwtan):

Did u mean the conjugate?

OpenStudy (anonymous):

after i multiplied both by the conjugate would i get that?

OpenStudy (anonymous):

yes no?

OpenStudy (cwtan):

I think no.

OpenStudy (anonymous):

i don't understand this prob. at all

OpenStudy (cwtan):

is sqrt -64 divide by (7-6i)-(2-2i) OR divide (7-6i) then whole thing minus (2-2i)?

OpenStudy (anonymous):

im sorry i still do not understand it

OpenStudy (cwtan):

\[\frac{\sqrt {-64}}{(7-6i)-(2-2i)}\] or \[\frac{\sqrt{-64}}{7-6i}-(2-2i)\] which one?

OpenStudy (anonymous):

the second one

OpenStudy (cwtan):

ok sqrt -64=8i 8i(7+6i)=56i-48--->(Note that i x i=-1) (7-6i)(7+6i)=49+36=85--->Using(a+b)(a-b)=a^2-b^2(also note that i^2=-1) so it is \[\frac{-48+56i}{85}\]

OpenStudy (cwtan):

after that u can do normal subtraction

OpenStudy (anonymous):

how?